There are two isotopes of an unknown element, X-19 and X-21. The abundance of X-19 is 12.01%.


What is the percentage of the X-21 isotope?


1
Expert's answer
2021-02-23T04:35:03-0500

Solution:

The abundance of X-19 + The abundance of X-21 = 100%

12.01% + The abundance of X-21 = 100%

Hence,

The abundance of X-21 = 100% - 12.01% = 87.99%

The abundance of X-21 = 87.99%


Answer: The percentage of the X-21 isotope is 87.99%.